OpenStudy (anonymous):

What is the sum of the angle measures of a 37-gon?

OpenStudy (anonymous):

35(180) a la 180(n-2)

OpenStudy (ash2326):

Sum of internal angles of a polygon with no. of sides n= (n-2)*180 for a 37 gon, n= 37 so sum of angles= (37-2)*180=35*180=6300

OpenStudy (anonymous):

Man, ash, way to show me up with a better answer.

OpenStudy (anonymous):

(180-360/37) * 37

OpenStudy (anonymous):

Don't use the "formula". Use the approach of triangles and angles. :)